There are 4 blue balls, 2 red balls, and 5 green balls in a bag. James removes one ball and does not replace it. James removes a second ball. What is the probability that both balls will be the same colour.

Total number of balls = 4 + 2 + 5 = 11 (this is our denominator for first draw)Numerator is the number of each colour. When we remove a ball the denominator and numerator for that colour decrease by 1. P(2 blue) = (4/11)(3/10) = (12/110), P(2 red) = (2/11)(1/10) = (2/110), P(2 green) = (5/11)(4/11) = (20/110)To find the total probability of picking two of the same colour we add these probabilities together then simplify.Total probability = 12/110 + 2/110 + 20/110 = 34/110 = 17/55

CW

Related Maths GCSE answers

All answers ▸

Solve the simultaneous equations (1) 2a - 5b = 11 (2) 3a + 2b = 7


What is (18)^1/2 -(8)^1/2 simplified?


Line A passes through point Q(2,3). Line B is parallel to line A and has the equation 2y-3x=4. What is the equation of line A?


Question from an Edexcel GCSE Maths Higher Paper (Nov 2018) - Solve the simultaneous equations: 5x + y = 21, x-3y = 9 (3 marks)